# FoodLMM: A Versatile Food Assistant using Large Multi-modal Model

### v1

@@ 8.1 Implementation Details
Table 12: Performance comparison on Food-101 [3]
@@
Table 13: Performance comparison on VIREO Food-172 [6]
@@
Table 14: Performance comparison on Recipe1M [79]
8.3 Ablation experiment of the ratio of one-to-zero samples.
One-to-Zero One-to-One/One-to-Many
$refer@1$ $refer@2$ refer@3 $refer@1$ $refer@2$ refer@3 refer@4 refer@5
Table 15: Ablation study of One-to-Zero referring data ratio.
8.4 Additional Results of Food VQA task
@@ 8.5 More Qualitative Examples
Figure 22 and Figure 23 present additional qualitative results from multi-turn dialogue tasks on nutrition-related topics. In Figure 24, we display additional qualitative examples from the reasoning segmentation task.
Figure 22: Qualitative examples of multi-round conversation.

### v2

@@ 8.2 Additional results of the referring segmentation task
One-to-Zero One-to-One/One-to-Many
$refer@1$ $refer@2$ refer@3 $refer@1$ $refer@2$ refer@3 refer@4 refer@5
Table 13: Ablation study of One-to-Zero referring data ratio.
8.3 Ablation experiment of the ratio of one-to-zero samples
8.4 Additional Results of Food VQA tasks
@@
Table 14: Performance comparison on Food-101 [3].
@@
Table 15: Performance comparison on VIREO Food-172 [8].
Method F1 score
DSDL(Resnet101) [101] 49.44
AFN+BFL(Resnet101) [56] 58.53
ML-GCN(Resnet101) [16] 55.32
SGTN(Resnet101) [86] 53.97
MGTN(Resnet101) [71] 55.08
FCI-MTL(Resnet101) [90] 59.56
DB-Focal(Resnet50) [95] 60.93
ASL(Resnet101) [77] 63.08
D-Mixup(Resnet101) [30] 63.16
CACLNet(Resnet101) [59] 65.71
FoodLMM FT 68.97
Table 16: Performance comparison on VIREO Food-172 [8].
@@
Table 17: Performance comparison on Recipe1M [81].
@@ 8.5 More Qualitative Examples
Figure 17 present additional qualitative results from multi-turn dialogue tasks on nutrition-related topics. In Figure 18, we display additional qualitative examples from the reasoning segmentation task.